the area of a right triangular rug is 12 squar feet. the length of the rug is 2 feet more then the width of the rug. what is the width od the right triangular rug?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

let the width of the rug be x feet
the length will be x+2
Area of triangle = 1/2 * side1 * side2 since it is right angled
12= 1/2 x(x+2)
24 = x^2 +2x
x^2 +2x-24 = 0
x^2+6x-4x-24 =0
x(x+6)-4(x+6)=0
(x+6)(x-4) =0
x=-6 or x=4
4feet is the width of the rug



Answer by stanbon(75887)

The area of a right triangular rug is 12 square feet. the length of the rug is 2 feet more then the width of the rug.
what is the width of the right triangular rug?
---
Let the width be "x".
Then the length is "x+2".
---
Equation:
Area = (1/2)*width*length
---
12 = (1/2)x(x+2)
24 = x^2 + 2x
x^2 + 2x - 24 = 0
Factor:
(x+6)(x-4) = 0
Positive solution:
x = 4 ft. (width)
x+2 = 6 ft. (length)
=========================
Cheers,
Stan H.
=========================